# On the Mechanisms of Collaborative Learning in VAE Recommenders

### v1

@@ Appendix E Theoretical development
  • The Kantorovich-Rubinstein dual form (rubinstein1970problem):
@@
  • The Donsker-Varadhan variational formula (donsker1975asymptotic): for any measurable $g$ with $\int e^{g}\,dp<\infty$,
@@
  • The Bobkov–Götze/Talagrand $T_{1}(C)$ inequality (bobkov1999exponential) is known to be equivalent to the following sub-Gaussian moment generating function (mgf) bound for Lipschitz functions: $p$ satisfies the $T_{1}(C)$ inequality with constant $C>0$, then for every $1$-Lipschitz $f$ and every $\lambda\in\mathbb{R}$,
@@
Additionally, assume the prior $p$ satisfies the Bobkov–Götze/Talagrand $T_{1}(C)$ inequality (bobkov1999exponential) with constant $C>0$ (e.g., Normal prior $p=\mathcal{N}(0,\sigma^{2}I)$):
@@
Recall the convex conjugate $A^{*}$ and Fenchel–Young inequality (fenchel1949conjugate; rockafellar1970convex):
@@
with equality when $y\in\partial A(\eta)$; in the regular (Legendre) case (rockafellar1970convex), $A$ is essentially smooth and strictly convex, so $y=\nabla A(\eta)$ is the unique equality condition and $A^{*}$ is strictly convex on its (convex) effective domain.
@@
Löwner-order contraction under commutation (HighamLin2013ShortCourse).

### v2

@@ Appendix E Theoretical development
  • The Kantorovich-Rubinstein dual form (Santambrogio, 2015):
@@
  • The Donsker-Varadhan variational formula (Donsker and Varadhan, 1975): for any measurable $g$ with $\int e^{g}\,dp<\infty$,
@@
  • The Bobkov–Götze/Talagrand $T_{1}(C)$ inequality (Bobkov and Götze, 1999) is known to be equivalent to the following sub-Gaussian moment generating function (mgf) bound for Lipschitz functions: $p$ satisfies the $T_{1}(C)$ inequality with constant $C>0$, then for every $1$-Lipschitz $f$ and every $\lambda\in\mathbb{R}$,
@@
Additionally, assume the prior $p$ satisfies the Bobkov–Götze/Talagrand $T_{1}(C)$ inequality (Bobkov and Götze, 1999) with constant $C>0$ (e.g., Normal prior $p=\mathcal{N}(0,\sigma^{2}I)$):
@@
Recall the convex conjugate $A^{*}$ and Fenchel–Young inequality (Fenchel, 1949; Rockafellar, 1970):
@@
with equality when $y\in\partial A(\eta)$; in the regular (Legendre) case (Rockafellar, 1970), $A$ is essentially smooth and strictly convex, so $y=\nabla A(\eta)$ is the unique equality condition and $A^{*}$ is strictly convex on its (convex) effective domain.
@@
Löwner-order contraction under commutation (Higham and Lin, 2013).